POSITION DESCRIPTION FOR THE SCGP ART PROGRAM DI-RECTOR/CURATOR

Description:

The Art Program Director/Curator position is part time in the Simons Center for Geometry and Physics and part time in the Department of Art at Stony Brook University. The Director/Curator will oversee the artistic activities of the Simons Center for Geometry and Physics, including the selection of art for the SCGP art gallery and other exhibition spaces in and around the building. The Art Program Director will also arrange art-and- science workshops, lectures, and outreach programs in addition to staged readings, concerts and other cultural events to benefit the members of the SCGP as well as the general University and Stony Brook communities. The incumbent will also be responsible for managing the program budget of $100,000 annually and will report regularly to the SCGP Arts Program Advisory Board. The Art Program Director will also teach one course a semester in Curatorial and/or Museum Studies in the Department of Art, alternating as needed with courses in the incumbents area of specialty, whether in Studio or History and Criticism.

Terms:

This is a full-time, five-year, Visiting Scholar position split between the SCGP and the Art Department with an annual stipend of a visiting Assistant Professor, around $70,000.

Qualifications:

Candidates must have an MFA degree in studio art or an MA/PhD in art history and criticism, and at least 5 years experience in art program administration and curation, outreach and events planning. Proven record of engagement in the sciences is highly preferred, as is experience with performance, installation art, and musical presentations.

Duties:

The Art Program Director, with advice from the Advisory Board, is responsible for producing public art and science events at the SCGP and elsewhere on campus, including the organization of exhibitions in the SCGP plaza-level Gallery and Penrose lobby exhibition spaces, art-and-science workshops, concerts, lectures and other academically oriented art/cultural activities. The Art Program Director will work with SCGP and campus staff members to ensure that all exhibits and other events follow University and State guidelines with regard to insurance, health and safety. Based on the incumbents area of expertise, the Art Program Director will teach (one per semester) undergraduate courses and/or graduate courses from the programs of the MFA in Studio Art or the PhD in Art History and Criticism. The incumbent may be responsible for supervising student projects and participating in critiques of studio projects, provided those activities do not conflict with responsibilities at the SCGP.

Special Notes:

With regard to budgetary oversight, the Art Program Director is responsible for the following:

• At the beginning of each Academic Year the Program Director shall prepare a proposed budget for the following year and present it for approval to the Art Program Advisory Board for its approval.

• The Program Director shall submit to the Board, at the end of each semester, a report on actual and adjusted planned expenses from this budget.

• The budget should largely be used for is intended to cover expenses associated with exhibitions, lectures, workshops, concerts and staged plays and/or readings.

• The Art Program Advisory Board must approve in advance all major expenditures, including equipment purchases.